# Steering Large Language Models with Register Analysis for Arbitrary Style Transfer

### v1

@@ 4.1.1 Authorship Imitation
We use LUAR  (Rivera-Soto et al., 2021), an authorship embedding model trained on MUD, to embed input, target and rewritten texts in the authorship space. We calculate the ”Away” and ”Towards” scores to represent the rewritten-input distance and the rewritten-output distance in the authorship embedding space respectively:
@@ Appendix B Target Construction
To avoid accidentally exposing the ”gold answer” to the rewriting systems, we select targets from the train split of the GYAFC dataset. Texts in GYAFC are of sentence-level and thus may be too short to contain enough linguistic information to inform the LLM. Hence, for each input text, we concatenate $K=16$ texts randomly selected from the target selection pool to form a paragraph-level target, following the practices of our MUD evaluation set construction for authorship imitation. Targets have the same domain and opposite formality with their corresponding input text. For example, for an formal input text in the ”EM” domain, 16 sentence-level targets are randomly selected from the EM-train-informal split of GYAFC to form a paragraph-level target to inform style during model’s inference time.
@@
To avoid accidentally exposing the ”gold answer” to the rewriting systems, we randomly select plain-language summary texts from the train split of the Cochrane dataset as target. Unlikely, texts in GYAFC are of paragraph-level, so each of them suffices in length to serve as a single target without the need of concatenation.
@@ Appendix D Implementation Details of Biber’s MDA Style Representation
The training and inference phrase of Biber’s MDA representation are shown in Fig 4. We train one Biber’s MDA representation per task (MUD, GYAFC, Cochrane). The training corpus for each task encompass a diverse range of styles (authorship for MUD, formality for GYAFC, and simplicity for Cochrane). The data splits constituting the training corpus for each task are specified as below:

### v2

@@ 4.1.1 Authorship Imitation
We use LUAR  (Rivera-Soto et al., 2021), an authorship embedding model trained on MUD, to embed input, target and rewritten texts in the authorship space. We calculate the ”Away” and ”Towards” scores to represent the rewritten-input distance and the rewritten-target distance in the authorship embedding space respectively:
@@ Appendix B Target Construction
To avoid accidentally exposing the ”gold answer” to the rewriting systems, we select targets from the train split of the GYAFC dataset. Texts in GYAFC are of sentence-level and thus may be too short to contain enough linguistic information to inform the LLM. Hence, for each input text, we concatenate $K=16$ texts randomly selected from the target selection pool to form a paragraph-level target exemplar, following our MUD evaluation set construction practices for authorship imitation. Targets have the same domain and opposite formality as their corresponding input text. For example, for an formal input text in the ”EM” domain, 16 sentence-level texts are randomly selected from the EM-train-informal split of GYAFC to form a paragraph-level target to inform style during model’s inference time.
@@
To avoid accidentally exposing the ”gold answer” to the rewriting systems, we randomly select plain-language summary texts from the train split of the Cochrane dataset as target. Unlike GYAFC, texts in Cochrane are of paragraph-level, so each of them suffices in length to serve as a single target without the need of concatenation.
@@ Appendix D Implementation Details of Biber’s MDA Style Representation
The train and inference phrase of Biber’s MDA representation are shown in Fig 4. We train one Biber’s MDA representation per task (MUD, GYAFC, Cochrane). The training corpus for each task encompasses a diverse range of styles for each task (authorship for MUD, formality for GYAFC, and simplicity for Cochrane). The data splits constituting the training corpus for each task are specified as follows: